Dear Richard dear all,
please find below a short summary of the activities of today on the mountain related to the AO system. Italian people working at LBTO were:
Arcetri: P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
ADS: P. Lazzarini
Total # of person: 8 W unit
1) The required mechanical machining of the AGW has been successfully done by the LBT staff
2) The AGW has been succesfully installed on the derotator of the FBGDX (please find attached the first picture of the LBT672 and AGW installed on the telescope). The definitive power and the cooling lines has been implemented to the AGW, while the fiber connection has been implemented temporarily waiting for the procurement of the long fiber.
3) A functionality test of the W unit has been successfully completed showing all the W subsystem working properly after the instalaltion on the telescope and at -2C of ambient temperature.
4) The "W alignment with rotator step1" has been performed. The measurement showed that the wobble is lower than 20 arcsec, but pointed out a tilt of the derotator of about 1.5arcmin for a given derotation range. This will be further investigated by A. Rakich with the laser tracker. AO control SW
1) Adsec Arbitrator has been modified to include TSS and telescope elevation control.
2) Meeting with N. Cushing, L. Fini for the schedule of the AOS installation on the telescope LBT672a
1) One defective actuator (on the external ring) has been identified and removed from the control loop.
2) The shell has been successfully tested at different elevation angles. The actuator current has been found in the expected range and no RIP event occurred.
3) The retroreflector optics group alignment has been checked in the clean room.
4) The installation of the laser tracker retroreflector optics on the LBT672 hub is now in progress.
Next update tomorrow
All the best
- 02 Apr 2010